THEORETICAL INVESTIGATION OF A NONADIABATIC MECHANISM FOR ALKYLPERNITRITE DE-COMPOSITION

\underline{JOHN M. HERBERT}, Department of Chemistry, University of California, Berkeley, CA 94720; ANNE B. McCOY, Department of Chemistry, The Ohio State University, Columbus, OH 43210; and JOHN F. STANTON, Department of Chemistry and Biochemistry, University of Texas, Austin, TX 78712.

We investigate a nonadiabatic two-state mechanism for reaction of alkylperoxyl radicals (ROO) with nitric oxide, an important step in the tropospheric ozone cycle. In our proposal, which is supported by high-level ab initio electronic structure calculations for R = CH_{3}, the pernitrite adduct ROONO propagates initially on a diabatic potential surface that correlates asymptotically with the Ã ^{2}B_{2} state of NO_{2}. This potential intersects a second diabatic potential corresponding to the stable nitrate species RNO_{2}, which correlates with ground-state (X̃ ^{2}A_{1}) NO_{2}. Near the intersection of these two surfaces, ROONO exhibits substantial diradical character. Diabatic potentials for the model system FONO have been developed based on CCSD(T)/cc-pVTZ electronic structure calculations. Here we report the results of classical and quantum dynamics simulations for FO + NO → FONO → FNO_{2}.